# Robust, fast, and efficient formation of stable tetratomic molecules from ultracold atoms via generalized stimulated Raman exact passage

### v3

@@ I Introduction
The successful creation of various ultracold diatomic molecules has stimulated growing interest in polyatomic molecules (UPMs) Rittenhouse and Sadeghpour (2010); Zeppenfeld et al. (2012); Kozyryev et al. (2017); Yang et al. (2022); Vilas et al. (2024); Chen et al. (2024). In contrast to diatomic molecules, UPMs offer additional complexity Carr et al. (2009); Küpper et al. (2009); Patterson et al. (2010), lending themselves to numerous interesting studies in, for instance, cold chemistry Karman et al. (2024), precision measurements Hutzler (2020) and the realization of exotic quantum phases Schmidt et al. (2022), to name a few. In order to tap their full potential, however, long-lived ensembles and full quantum state control serve as prerequisites Langen et al. (2024). Thus, the efficient conversion of ultracold atoms into stable UPMs constitutes an essential requirement, which hitherto remains challenging.
@@
In the ongoing experiments, ultracold molecules are first prepared by association of pre-cooled atoms Chin et al. (2010), either via photo- or the magneto-association Dulieu and Gabbanini (2009); Chandler (2010); Quéméner and Julienne (2012); Zhao and Pan (2022). Subsequently, the formed molecules should be immediately transferred to absolute ground states, it is commonly achieved via the well-known Stimulated Raman adiabatic passage (STIRAP) Drummond et al. (2002); Lang et al. (2008) and its variants, such as multi-state STIRAP Danzl et al. (2010); Mark et al. (2009); Kuznetsova et al. (2008); Jing et al. (2008); Qian et al. (2010); Devolder et al. (2021). In principle, by manipulating the system to evolve along the adiabatic dark state, an efficient and robust state transfer can be obtained Bergmann et al. (1998); Král et al. (2007); Vitanov et al. (2017). Until now, STIRAP has been demonstrated successfully Bergmann et al. (2019), however, its application has been limited to the realm of diatomic molecules Winkler et al. (2007); Lang et al. (2008); Danzl et al. (2008); Aikawa et al. (2010); Ospelkaus et al. (2008); Feng et al. (2009); Liu et al. (2019a); Duda et al. (2023); Zamarski et al. (2025). Directly applying it to UPMs preparation remains challenging. To address this problem, researchers have developed STIRAP variants, for instance, Feshbach/Efimov resonance-assisted STIRAP, to support theoretical investigations into UPMs preparation Jing et al. (2007); Meng et al. (2009); Li and Peng (2011); Dou et al. (2012). Nevertheless, the successful application of such techniques is limited by the adiabatic criterion Zhou et al. (2010), which renders the systems vulnerable to environmental factors, e.g., collisions or decay to outside states.
@@
In order to relax limitations such as adiabaticity related to the STIRAP-based scheme, a technique which is called shortcut-to-adiabaticity (STA) was proposed Chen et al. (2010). By leveraging the STA, the adiabatic process can be expedited while ensuring that the outcomes remain consistent. One particularly successful method for designing such fast routes is counterdiabatic driving Demirplak and Rice (2005, 2008); Berry (2009); Finžgar et al. (2025); Morawetz and Polkovnikov (2025); Dengis et al. (2025a, b), which allows one to construct a modification of an original Hamiltonian to eliminate unwanted nonadiabatic couplings Guéry-Odelin et al. (2019). Currently, STA is widely used in theory to realize fast and robust ultracold molecule formation Zhu and Chen (2021); Zhang and Dou (2021); Cao et al. (2023) and control the selective vibrational population transfer of molecules Masuda and Rice (2015a, b). In a recent study, Zhu etet alalZhu and Chen (2021) proposed a fast-forward scaling-assisted STIRAP method to study the conversion from atoms to diatomic molecules. In another study, Vitanov developed a shortcut multi-state STIRAP technique Vitanov (2020), which may provide an alternative ways for forming stable ground-state ultracold molecules. Compared to the traditional three-level system, the employment of multi-level ones exhibits unique advantages; to illustrate, this strategy proves instrumental in resolving the challenge of weak Franck-Condon factors (FCFs) during the process of both atom-to-molecule conversion Jaksch et al. (2002); Qian et al. (2010) and ground-state molecular transfer Danzl et al. (2010); Kuznetsova et al. (2008). Despite these advances, the major problem is that most prior studies usually require introducing additional couplings for eliminating non-adiabatic effects, which may pose additional experimental challenges. On the other hand, in 2017, Dorier etet alalDorier et al. (2017) suggested an alternative method, namely nonlinear stimulated Raman exact passage (STIREP), for the efficient atom-to-molecule conversion. This method conforms to the spirit of STA, because the control fields are determined by the specific dynamics selected to achieve exact state transfer Laforgue et al. (2022); Dridi et al. (2024). This trick has been extended to the formation of triatomic molecules Cao (2024). Currently, an important question in this realm is whether it is possible to extend this successful tool to form larger polyatomic molecular complexes Dou et al. (2013); Shammout et al. (2025); Quéméner et al. (2023); Cao et al. (2024); Gacesa et al. (2021); Gu et al. (2025), and generalize it to transfer of molecules into deeply-bound states Koch and Moszyński (2008); Danzl et al. (2010); Kuznetsova et al. (2008); de Lima (2024); Mukherjee and Tomza (2025).
@@ II.1 Model and Method
Figure 1: (Color online) Schematic showing the coherent conversion from ultracold atoms to tetratomic molecules.
@@
Starting from a Bose condensate of atoms, we apply the field $\Omega_{1}^{\prime}$ to associate free atoms $|a\rangle$ into triatomic molecules $|m\rangle$; the other field $\Omega_{2}^{\prime}$, in turn, is responsible for associating pairs composed of an atom and a triatomic molecule $|m\rangle$ into tetratomic molecules $|t\rangle$. For the applied lasers, their frequencies are expressed using the single-photon and two-photon detunings $\delta$ and $\Delta$, respectively. We remark that the formation of triatomic molecules via the photoassociation of free atoms has been recently reported Schnabel et al. (2021), whereas the formation of tetratomic molecules through the coupling of an atom-triatomic molecule pair has also been investigated Jing and Jiang (2008). Our aim is to find an exact solution for controlling fast and efficient conversion from ultracold atoms to tetratomic molecules by generalizing existing N-STIREP protocols.
@@
The Hamiltonian of this system, in the interaction picture with $\hbar=1$, takes the form of
@@
$\begin{aligned} \hat{\mathcal{H}}\!=&-\int\!d\mathbf{r}\left\{\sum_{ij}\chi_{ij}^{\prime}\hat{\psi}_{i}^{\dagger}(\mathbf{r})\hat{\psi}_{j}^{\dagger}(\mathbf{r})\hat{\psi}_{j}(\mathbf{r})\hat{\psi}_{i}(\mathbf{r})\!+\!\delta\hat{\psi}_{\mathrm{m}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{m}}(\mathbf{r})\right.\\ &\!+\!\Omega_{1}^{\prime}\{\hat{\psi}_{\mathrm{m}}^{\dagger}(\mathbf{r})[\hat{\psi}_{\mathrm{a}}(\mathbf{r})]^{3}+\mathrm{H.c.}\}\!+\!(\Delta\!+\!\delta)\hat{\psi}_{\mathrm{t}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{t}}(\mathbf{r})\\ &\!-\!\Omega_{2}^{\prime}[\hat{\psi}_{\mathrm{t}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{m}}(\mathbf{r})\hat{\psi}_{\mathrm{a}}(\mathbf{r})+\mathrm{H.c.}]\Bigg\}.\end{aligned}$ (1)
@@
where $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ and $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ are, respectively, the bosonic creation and annihilation operators for state $|k\rangle(k\!=\!a,m,t)$. The indices $i,j=a,m,t$ represent the atomic, triatomic molecular, and tetratomic molecular states, respectively. By adopting the mean-field approach $\hat{\psi}_{i}\rightarrow\sqrt{n}{\psi}_{i}$ ($n$ denotes the initial atomic density) Ling et al. (2004), the Heisenberg motional equation can be described as
$\begin{aligned} i\dot{\psi}_{a}\!&=\!-2K_{a}\psi_{a}\!-\!3\Omega_{1}{\psi_{m}\psi^{\ast}_{a}}^{2}\!+\!\Omega_{2}\psi_{t}{\psi^{*}_{m}},\\ i\dot{\psi}_{m}\!&=\!-2K_{m}\psi_{m}\!-\!(\delta\!+\!i\gamma_{m})\psi_{m}\!-\!\Omega_{1}\psi_{a}^{3}\!+\!\Omega_{2}\psi_{t}{\psi^{\ast}_{a}},\\ i\dot{\psi}_{t}\!&=\!-2K_{t}\psi_{t}\!-\!(\Delta\!+\!\delta\!+\!i\gamma_{t})\psi_{t}\!+\!\Omega_{2}\psi_{m}\psi_{a}.\end{aligned}$ (2)
@@
In the above equations, $\chi_{ij}\!=\!n\chi_{ij}^{\prime},\Omega_{1}\!=\!n\Omega_{1}^{\prime},$ and $\Omega_{2}\!=\!\sqrt{n}\Omega_{2}^{\prime}$ denote the renormalized quantities. $\gamma_{m}$ and $\gamma_{t}$ are introduced to phenomenologically describe the losses of the corresponding states to other undetected states. Nevertheless, we temporarily ignore these losses in the theoretical analysis to clarify the underlying physics. In this case, we have $|\psi_{\mathrm{a}}|^{2}\!+\!3|\psi_{\mathrm{m}}|^{2}\!+\!4|\psi_{\mathrm{t}}|^{2}\!=\!1$. The third-order nonlinearity term $K_{i}\!=\!\sum_{j}\chi_{ij}\left|\psi_{j}\right|^{2}$ characterizes elastic collisions between particles, which gives rise to the dynamical instability effect Itin and Watanabe (2007); Guérin et al. (2013).
@@
In principle, the effects of third-order nonlinearity can be eliminated by employing resonance locked inverse engineering Dorier et al. (2017); Zhu et al. (2023); Cao et al. (2023). To this end, a general parametrization with the phase $\alpha$ should be introduced, i.e., $\psi_{\mathrm{a}}\!\mapsto\!\psi_{\mathrm{a}}e^{i\alpha},\psi_{\mathrm{m}}\!\mapsto\!\psi_{\mathrm{m}}e^{\mathrm{3}i\alpha},\psi_{\mathrm{t}}\!\mapsto\!\psi_{\mathrm{t}}e^{\mathrm{4}i\alpha}$. From this, the condition for the resonance locking can be derived as
@@
$\begin{aligned} \delta=6K_{\mathrm{a}}\!-\!2K_{\mathrm{m}},\quad\Delta=2(K_{\mathrm{a}}\!+\!K_{\mathrm{m}}\!-\!K_{\mathrm{t}}).\end{aligned}$ (3)
@@
$\begin{aligned} i\dot{\psi}_{\mathrm{a}}\!&=\!-\Omega_{1}{\psi_{\mathrm{m}}\psi^{\ast}_{\mathrm{a}}}^{2}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{*}_{\mathrm{m}}},\\ i\dot{\psi}_{\mathrm{m}}\!&=\!-\Omega_{1}\psi_{\mathrm{a}}^{3}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{\ast}_{\mathrm{a}}},\\ i\dot{\psi}_{\mathrm{t}}\!&=\!\Omega_{2}\psi_{\mathrm{m}}\psi_{\mathrm{a}}.\end{aligned}$ (4)
Now two optical fields $\Omega_{1,2}$ are resonant with the transitions $|\mathrm{a}\rangle\leftrightarrow|\mathrm{m}\rangle$ and $|\mathrm{m}\rangle\leftrightarrow|\mathrm{t}\rangle$, respectively.
We further note that, in our model, the couplings of both fields are nonlinear, which differs from Ref. Dorier et al. (2017), where only the coupling of the first field is nonlinear. In what follows, we demonstrate how the generalized STIREP adapts to the system to compensate for the nonlinearities induced by these couplings. To this end, we introduce two dynamical angles $\theta$ and $\phi$ to parameterize the state vector:
$\begin{aligned} &\psi_{\mathrm{a}}\!=\!\cos\phi\cos\theta,\\ &\psi_{\mathrm{m}}\!=\!-i\frac{\sin\phi}{\sqrt{3}},\\ &\psi_{\mathrm{t}}\!=\!-\frac{\cos\phi\sin\theta}{2}.\end{aligned}$ (5)
@@
$\begin{aligned} \Omega_{1}&\!=\!-\frac{\sqrt{3}}{12}\left(\frac{1\!+\!3\cos^{2}\theta}{\cos^{3}\theta\cos^{2}\phi}\dot{\phi}\!+\!\frac{10\tan\theta\dot{\theta}}{\sin 2\phi\cos\theta}\right),\\ \Omega_{2}&\!=\!\frac{\sqrt{3}}{2}\left(\frac{\dot{\theta}}{\sin\phi}\!-\!\frac{\dot{\phi}\tan\theta}{\cos\phi}\right).\end{aligned}$ (6)
To ensure the desired atom-to-tetratomic molecular conversion, we apply boundary conditions: $0\!\leftarrow\!\phi\!\rightarrow\!0$ and $0\!\leftarrow\!\theta\!\rightarrow\!\pi/2$, where the arrows to the left and right signify the limits when $t\!\rightarrow\!t_{i}$ and $t\!\rightarrow\!t_{f}$, respectively. The controlled parameters $\theta$ and $\phi$ are chosen as
$\begin{aligned} \theta(t)&=\frac{\pi}{4}\eta\left(\mathrm{tanh}\left(\frac{t}{T}\right)\!+\!1\right),\\ \phi(t)&=\frac{4\epsilon\sqrt{\theta(t)\left[\frac{\pi}{2}\!-\!\theta(t)\right]}}{\pi},\end{aligned}$ (7)
where $\epsilon$ and $\eta$ are constants. $\epsilon\gtrsim 0$ enables the regulation of the transient population in the intermediate state, while $\eta\gtrsim 1$ governs the final conversion efficiency. It should be noted that $\eta=1$ is prohibited in nonlinear cases, whereas no such restriction exists in linear cases.

### v4

@@ I Introduction
The successful creation of various ultracold diatomic molecules has stimulated growing interest in polyatomic molecules (UPMs) Zeppenfeld et al. (2012); Kozyryev et al. (2017); Yang et al. (2022); Quéméner et al. (2023); Vilas et al. (2024); Chen et al. (2024) In contrast to diatomic molecules, UPMs offer additional complexity Carr et al. (2009); Küpper et al. (2009); Patterson et al. (2010), lending themselves to numerous interesting studies in, for instance, cold chemistry Karman et al. (2024), precision measurements Hutzler (2020) and the realization of exotic quantum phases Schmidt et al. (2022), to name a few. In order to tap their full potential, however, long-lived ensembles and full quantum state control serve as prerequisites Langen et al. (2024). Thus, the efficient conversion of ultracold atoms into stable UPMs constitutes an essential requirement, which hitherto remains challenging.
@@
In the ongoing experiments, ultracold molecules are first prepared by association of pre-cooled atoms Chin et al. (2010), either via magneto or the photo-association (PA) Dulieu and Gabbanini (2009); Chandler (2010); Quéméner and Julienne (2012); Zhao and Pan (2022). Subsequently, the formed molecules should be immediately transferred to absolute ground states, it is commonly achieved via the well-known Stimulated Raman adiabatic passage (STIRAP) Drummond et al. (2002); Lang et al. (2008) and its variants, such as multi-state STIRAP Danzl et al. (2010); Mark et al. (2009); Kuznetsova et al. (2008); Jing et al. (2008); Qian et al. (2010); Devolder et al. (2021). In principle, by manipulating the system to evolve along the adiabatic dark state, an efficient and robust state transfer can be obtained Bergmann et al. (1998); Král et al. (2007); Vitanov et al. (2017). Until now, STIRAP has been demonstrated successfully Bergmann et al. (2019), however, its application has been limited to the realm of diatomic molecules Winkler et al. (2007); Lang et al. (2008); Danzl et al. (2008); Aikawa et al. (2010); Ospelkaus et al. (2008); Feng et al. (2009); Liu et al. (2019a); Duda et al. (2023); Zamarski et al. (2025). Directly applying it to UPMs preparation remains challenging. On the other hand, researchers have developed some generalized versions of STIRAP, for instance, Feshbach/Efimov resonance-assisted STIRAP, to support theoretical investigations into UPMs preparation Jing et al. (2007); Meng et al. (2009); Li and Peng (2011); Dou et al. (2012). Nevertheless, the successful application of such techniques is limited by the adiabatic criterion, which renders the systems vulnerable to environmental factors, e.g., collisions or decay to outside states.
@@
In order to relax limitations such as adiabaticity related to the STIRAP-based scheme, a technique which is called shortcut-to-adiabaticity (STA) was proposed Chen et al. (2010). By leveraging the STA, the adiabatic process can be expedited while ensuring that the outcomes remain consistent. One especially successful method for designing such fast routes is counterdiabatic driving Demirplak and Rice (2005, 2008); Berry (2009); Finžgar et al. (2025); Morawetz and Polkovnikov (2025); Dengis et al. (2025), which allows one to construct a modification of an original Hamiltonian to eliminate unwanted nonadiabatic couplings Guéry-Odelin et al. (2019). Currently, STA is widely used in theory to realize fast and robust ultracold molecule formation Zhu and Chen (2021); Zhang and Dou (2021); Cao et al. (2023) and control the selective vibrational population transfer of molecules Masuda and Rice (2015a, b). In a recent study, Zhu etet alalZhu and Chen (2021) proposed a fast-forward scaling-assisted STIRAP method to study the conversion from atoms to diatomic molecules. In another study, Vitanov developed a shortcut multi-state STIRAP technique Vitanov (2020), which may provide an alternative ways for forming stable ground-state ultracold molecules. Very recently, a similar study has been reported Han et al. (2024). Compared to the traditional three-level system, the employment of multi-level ones exhibits unique advantages; to illustrate, this strategy proves instrumental in resolving the challenge of weak Franck-Condon factors (FCFs) during the process of both atom-to-molecule conversion Jaksch et al. (2002); Qian et al. (2010) and ground-state molecular transfer Danzl et al. (2010); Kuznetsova et al. (2008). Despite these advances, the major problem is that most prior studies usually require introducing additional couplings for eliminating non-adiabatic effects, which may pose additional experimental challenges. On the other hand, in 2017, Dorier etet alalDorier et al. (2017) suggested an alternative method, namely nonlinear stimulated Raman exact passage (STIREP), for the efficient atom-to-molecule conversion. This method conforms to the spirit of STA, because the control fields are determined by the specific dynamics selected to achieve exact state transfer Laforgue et al. (2022); Dridi et al. (2024). This trick has been extended to the formation of triatomic molecules Cao (2024). Currently, an important question in this realm is whether it is possible to extend this successful tool to form larger polyatomic molecular complexes, Dou et al. (2013); Shammout et al. (2025); Quéméner et al. (2023); Cao et al. (2024); Gacesa et al. (2021); Gu et al. (2025) and generalize it to transfer of molecules into deeply-bound ground states. Koch and Moszyński (2008); Danzl et al. (2010); Kuznetsova et al. (2008); de Lima (2024)
@@ II.1 Model and Method
Figure 1: (Color online) Schematic showing the coherent conversion from ultracold atoms to tetratomic molecules Jones et al. (2006).
@@
Starting from a Bose condensate of atoms, we apply the field $\Omega_{1}^{\prime}(t)$ to associate free atoms $|a\rangle$ into triatomic molecules $|m\rangle$; the other field $\Omega_{2}^{\prime}$(t), in turn, is responsible for associating pairs composed of an atom and a triatomic molecule $|m\rangle$ into tetratomic molecules $|t\rangle$. For the applied lasers, their frequencies are expressed using the single-photon and two-photon detunings $\delta(t)$ and $\Delta(t)$, respectively. We remark that the formation of triatomic molecules via the PA of free atoms has been recently reported Schnabel et al. (2021), whereas the formation of tetratomic molecules through the coupling of an atom-triatomic molecule pair has also been investigated Jing and Jiang (2008). In addition, Jones etet alalJones et al. (2006) pointed out that a sufficiently dense atomic sample facilitates the combination of three atoms via PA to form triatomic molecules.
Drawing on the above theoretical advances and the well-established strategy for the stepwise coherent assembly of UPMs Jing et al. (2007), our aim is to generalize the existing N-STIREP protocol, with the goal of enabling fast and efficient conversion of atoms to tetratomic molecules, while ensuring the occupation of the intermediate triatomic molecule remains negligible.
@@
The Hamiltonian of this system, in the interaction picture with $\hbar=1$, takes the form of Jing and Jiang (2008); Dou et al. (2012)
@@
$\begin{aligned} \hat{\mathcal{H}}\!=&-\int\!d\mathbf{r}\left\{\sum_{ij}\chi_{ij}^{\prime}\hat{\psi}_{i}^{\dagger}(\mathbf{r})\hat{\psi}_{j}^{\dagger}(\mathbf{r})\hat{\psi}_{j}(\mathbf{r})\hat{\psi}_{i}(\mathbf{r})\!+\!\delta\hat{\psi}_{\mathrm{m}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{m}}(\mathbf{r})\right.\\ &\!+\!\Omega_{1}^{\prime}(t)\{\hat{\psi}_{\mathrm{m}}^{\dagger}(\mathbf{r})[\hat{\psi}_{\mathrm{a}}(\mathbf{r})]^{3}+\mathrm{H.c.}\}\!+\![\Delta(t)\!+\!\delta(t)]\hat{\psi}_{\mathrm{t}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{t}}(\mathbf{r})\\ &\!-\!\Omega_{2}^{\prime}(t)[\hat{\psi}_{\mathrm{t}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{m}}(\mathbf{r})\hat{\psi}_{\mathrm{a}}(\mathbf{r})+\mathrm{H.c.}]\Bigg\},\end{aligned}$ (1)
@@
where $\hat{\psi}_{i}^{\dagger}(\mathbf{r})$ and $\hat{\psi}_{i}{(\mathbf{r})}$ are, respectively, the bosonic creation and annihilation operators for state $|i\rangle(i\!=\!\mathrm{a,m,t})$. In the present analysis, we focus on a uniform system, and for this reason, the usual kinetic term has been omitted Ling et al. (2004). For systems that contain a sufficiently large number of particles, the mean-field method holds validity Parkins and Walls (1998). By adopting this method, we use the substitution $\hat{\psi}_{i}\rightarrow\sqrt{n}{\psi}_{i}$ ($n$ denotes the initial atomic density), and the Heisenberg motional equation can be described as
$\begin{aligned} i\dot{\psi}_{\mathrm{a}}\!&=\!-2K_{\mathrm{a}}(t)\psi_{\mathrm{a}}\!-\!3\Omega_{1}{\psi_{\mathrm{m}}\psi^{\ast}_{\mathrm{a}}}^{2}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{*}_{\mathrm{m}}},\\ i\dot{\psi}_{\mathrm{m}}\!&=\!-2K_{\mathrm{m}}(t)\psi_{\mathrm{m}}\!-\!(\delta\!+\!i\gamma_{\mathrm{m}})\psi_{\mathrm{m}}\!-\!\Omega_{1}\psi_{\mathrm{a}}^{3}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{\ast}_{\mathrm{a}}},\\ i\dot{\psi}_{\mathrm{t}}\!&=\!-2K_{\mathrm{t}}(t)\psi_{\mathrm{t}}(t)\!-\!(\Delta\!+\!\delta\!+\!i\gamma_{\mathrm{t}})\psi_{\mathrm{t}}\!+\!\Omega_{2}\psi_{\mathrm{m}}\psi_{\mathrm{a}}.\end{aligned}$ (2)
@@
In the above equations, $\chi_{ij}\!=\!n\chi_{ij}^{\prime},\Omega_{1}\!=\!n\Omega_{1}^{\prime},$ and $\Omega_{2}\!=\!\sqrt{n}\Omega_{2}^{\prime}$ denote the renormalized quantities. The decay rates $\gamma_{m}$ and $\gamma_{t}$ are introduced to phenomenologically describe the losses of the corresponding states caused by factors such as inelastic collisions and laser-induced dissociation, under the assumption that these decays dominate all other loss mechanisms. The time-dependent third-order nonlinearity term $K_{i}(t)\!=\!\sum_{j}\chi_{ij}\left|\psi_{j}\right|^{2}$ characterizes elastic collisions between particles, and gives rise to the dynamical instabilities Itin and Watanabe (2007); Guérin et al. (2013).
@@
In principle, The effects of third-order nonlinearity can be eliminated by resonance locked inverse engineering Dorier et al. (2017); Zhu et al. (2023). To this end, a general parametrization with the time-dependent phase $\alpha$ is introduced, i.e.,
$\begin{aligned} \psi_{\mathrm{a}}\!&\mapsto\!\psi_{\mathrm{a}}e^{i\alpha(t)},\\ \psi_{\mathrm{m}}\!&\mapsto\!\psi_{\mathrm{m}}e^{\mathrm{3}i\alpha(t)},\\ \psi_{\mathrm{t}}\!&\mapsto\!\psi_{\mathrm{t}}e^{\mathrm{4}i\alpha(t)}.\end{aligned}$ (3)
In this way, the above equation can be expressed as
@@
$\begin{aligned} i\dot{\psi}_{\mathrm{a}}\!&=\!-[2K_{\mathrm{a}}(t)\!-\!\dot{\alpha}]\psi_{\mathrm{a}}\!-\!3\Omega_{1}{\psi_{\mathrm{m}}\psi^{\ast}_{\mathrm{a}}}^{2}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{*}_{\mathrm{m}}},\\ i\dot{\psi}_{\mathrm{m}}\!&=\!-\left[2K_{\mathrm{m}}(t)\!-\!3\dot{\alpha}\!+\!\delta\right]\psi_{\mathrm{m}}\!-\!i\gamma_{\mathrm{m}}\psi_{\mathrm{m}}\!-\!\Omega_{1}\psi_{\mathrm{a}}^{3}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{\ast}_{\mathrm{a}}},\\ i\dot{\psi}_{\mathrm{t}}\!&=\!-[2K_{\mathrm{t}}(t)\!-\!4\dot{\alpha}\!+\!\Delta\!+\!\delta]\psi_{\mathrm{t}}(t)\!-\!i\gamma_{\mathrm{t}}\psi_{\mathrm{t}}\!+\!\Omega_{2}\psi_{\mathrm{m}}\psi_{\mathrm{a}}.\end{aligned}$ (4)
With the substitution $\dot{\alpha}=2K_{\mathrm{a}}(t)$ implemented, the condition for resonance locking can be derived as
$\begin{aligned} \delta&=6K_{\mathrm{a}}(t)\!-\!2K_{\mathrm{m}}(t),\\ \Delta&=2[K_{\mathrm{a}}(t)\!+\!K_{\mathrm{m}}(t)\!-\!K_{\mathrm{t}}(t)].\end{aligned}$ (5)
@@
$\begin{aligned} i\dot{\psi}_{\mathrm{a}}\!&=\!-\Omega_{1}{\psi_{\mathrm{m}}\psi^{\ast}_{\mathrm{a}}}^{2}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{*}_{\mathrm{m}}},\\ i\dot{\psi}_{\mathrm{m}}\!&=\!-\!i\gamma_{\mathrm{m}}\psi_{\mathrm{m}}\!-\Omega_{1}\psi_{\mathrm{a}}^{3}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{\ast}_{\mathrm{a}}},\\ i\dot{\psi}_{\mathrm{t}}\!&=\!-\!i\gamma_{\mathrm{t}}\psi_{\mathrm{t}}+\!\Omega_{2}\psi_{\mathrm{m}}\psi_{\mathrm{a}}.\end{aligned}$ (6)
Now two optical fields $\Omega_{1,2}$ are resonant with the transitions $|\mathrm{a}\rangle\leftrightarrow|\mathrm{m}\rangle$ and $|\mathrm{m}\rangle\leftrightarrow|\mathrm{t}\rangle$, respectively. The detailed derivation from Eq. (2) to (6) is presented in the Appendix A.
We further note that, in our model (6), the couplings of both fields are nonlinear, which differs from Ref. Dorier et al. (2017), where only the coupling of the first field is nonlinear. In what follows, we demonstrate how the generalized STIREP adapts to the system to compensate for the nonlinearities induced by these couplings. Consistent with previous theoretical treatments, we ignore all decay rates in Eq. (6) to clarify the underlying physics. For the case with the absence of losses ($\gamma_{m,t}\!=\!0$), we have $|\psi_{\mathrm{a}}|^{2}\!+\!3|\psi_{\mathrm{m}}|^{2}\!+\!4|\psi_{\mathrm{t}}|^{2}\!=\!1$. On this basis, we introduce two time-dependent angles $\theta(t)$ and $\phi(t)$ to parameterize the state vector:
$\begin{aligned} &\psi_{\mathrm{a}}\!=\!\cos\phi(t)\cos\theta(t),\\ &\psi_{\mathrm{m}}\!=\!-i\frac{\sin\phi(t)}{\sqrt{3}},\\ &\psi_{\mathrm{t}}\!=\!-\frac{\cos\phi(t)\sin\theta(t)}{2}.\end{aligned}$ (7)
@@
$\begin{split}\Omega_{1}(t)&\!=\!-\frac{\sqrt{3}}{12}\biggl(\frac{1\!+\!3\cos^{2}\theta(t)}{\cos^{3}\theta(t)\cos^{2}\phi(t)}\dot{\phi}(t)\\ &\qquad+\frac{10\tan\theta(t)\dot{\theta}(t)}{\sin 2\phi(t)\cos\theta(t)}\biggr),\\ \Omega_{2}(t)&\!=\!\frac{\sqrt{3}}{2}\biggl(\frac{\dot{\theta}(t)}{\sin\phi(t)}\!-\!\frac{\dot{\phi}(t)\tan\theta(t)}{\cos\phi(t)}\biggl).\end{split}$ (8)
To ensure the desired atom-to-tetratomic molecular conversion, we apply boundary conditions: $0\!\leftarrow\!\phi\!\rightarrow\!0$ and $0\!\leftarrow\!\theta\!\rightarrow\!\pi/2$, where the arrows to the left and right signify the limits when $t\!\rightarrow\!t_{i}$ and $t\!\rightarrow\!t_{f}$, respectively. The time-dependent angles $\theta(t)$ and $\phi(t)$ are chosen as
$\begin{aligned} \theta(t)&=\frac{\pi}{4}\eta\left[\mathrm{tanh}\left(\frac{t}{T}\right)\!+\!1\right],\\ \phi(t)&=\frac{4\epsilon\sqrt{\theta(t)\left[\frac{\pi}{2}\!-\!\theta(t)\right]}}{\pi},\end{aligned}$ (9)
where $\epsilon$ and $\eta$ are constants, $T$ serves as a time for normalization. $\epsilon\gtrsim 0$ enables the regulation of the transient population in the intermediate state, while $\eta\lesssim 1$ governs the final conversion efficiency. It should be noted that $\eta=1$ is prohibited in nonlinear cases, whereas no such restriction exists in linear cases. Once $\theta(t)$ and $\phi(t)$ are determined, the Rabi frequencies can be derived by substituting them into Eq. (8).